Job Costs Using a Plantwide Overhead Rate Naranjo Company designs industrial prototypes for outside companies. Budgeted overh3. What if the expected direct labor rate at the beginning of the year was $28 instead of $35? What would the overhead rate bJob Costs Using a Plantwide Overhead Rate Naranjo Company designs industrial prototypes for outside companies. Budgeted overhead for the year was $280,000, and budgeted direct labor hours were 16,000. The average wage rate for direct labor is expected to be $35 per hour. During June, Naranjo Company worked on four jobs. Data relating to these four jobs follow: Job 39 Job 40 Job 41 Job 42 Beginning balance $22,700 $35,500 $15,800 $1,700 Materials requisitioned 20,100 21,000 10,300 15,600 Direct labor cost 11,200 18,100 4,950 6,500 Overhead is assigned as a percentage of direct labor cost. During June, Jobs 39 and 40 were completed; Job 39 was sold at 115 percent of cost. (Naranjo had originally developed Job 40 to order for a customer; however, that customer was near bankruptcy and the chance of Naranjo being paid was growing dimmer. Naranjo decided to hold Job 40 in inventory while the customer worked out its financial difficulties. Job 40 is the only job in Finished Goods Inventory.) Jobs 41 and 42 remain unfinished at the end of the month.
Answer #1
1 Overhead rate=Budgeted overhead/Budgeted direct labor cost
Budgeted direct labor cost=Budgeted direct labor hours*Average wage rate=16000*35=$ 560000
Overhead rate=280000/560000=0.50=50% of direct labor cost
2 Naranjo company
Job-order cost sheets
Job 39 Job 40 Job 41 Job 42
Balance,june 1 22700 35500 15800 1700
Materials requisitioned 20100 21000 10300 15600
Direct labor cost a 11200 18100 4950 6500
Overhead applied a*50% 5600 9050 2475 3250
Total 59600 83650 33525 27050
3 New Budgeted direct labor cost=Budgeted direct labor hours*Average wage rate=16000*28=$ 448000
New overhead rate=Budgeted overhead/Budgeted direct labor cost=280000/448000=0.625=62.5% of direct labor cost
Cost of the jobs would be increased since overhead rate has increased from 50% to 62.5%